OpenStudy (anonymous):

PLEASE HELP…CALCULUS Find the slope of the graph of the function at the given point. Use the derivative feature of a graphing utility to confirm you results. f(x)=3-3/5t (3/5,2)

OpenStudy (anonymous):

A graphing calaculator

OpenStudy (anonymous):

could you explain to me completely?

OpenStudy (anonymous):

this is the function: \(\large f(t)=3-\frac{3}{5t} \) ????

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

ok... so take the derivative.... and evaluate that at \(\large x=\frac{3}{2} \)

OpenStudy (anonymous):

@dpaInc x=3/5 not 3/2

OpenStudy (anonymous):

\(\large f(t)=3-\frac{3}{5t}=3-\frac{3}{5}t^{-1} \) \(\large f'(t)=\frac{3}{5}t^{-2} \) \(\large \frac{3}{5}t^{-2}|_{x={3/5}}=\frac{3}{5}\cdot \frac{25}{9}=\frac{5}{3} \) oohh... thanks @ipm1988 .... i read your post before i entered this... :)

OpenStudy (anonymous):

@dpaInc no problem :)

OpenStudy (anonymous):

now... the problem states confirm with your graphing utility.... what calculator do you use?

OpenStudy (anonymous):

TI-73

OpenStudy (anonymous):

ok.... enter the function into your calculator as " \(\large y_1= \) "

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

graph it .... now press "calc" --> "dy/dx" -->"3/5" your calculator will give you "x=1.66667"

OpenStudy (anonymous):

yes it did

OpenStudy (anonymous):

isn't there a way to do it without a calculator though

OpenStudy (anonymous):

which is the calculator's approximation for 5/3... ok....????

OpenStudy (anonymous):

yeah.... it's in my third post here.... @ipm1988 also showed you in his docx file....

OpenStudy (anonymous):

Ok thanks. for some reason I can't open @ipm1988s up but thanks for your help I think I got it.

OpenStudy (anonymous):

did you understand how i got the derivative?

OpenStudy (anonymous):

i just rewrote the function so i can use the power rule....

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I understood

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

ok... so the thing we did with the calculator just confirms that our answer is correct...

OpenStudy (anonymous):

yw... :)